# Error Broadcast and Decorrelation as a Potential Artificial and Natural Learning Mechanism

### v2

@@ 4 Numerical experiments
In this section, we evaluate the performance of the proposed Error Broadcast and Decorrelation (EBD) approach on two benchmark datasets: MNIST [44] and CIFAR-10 [45]. For experiments involving MLP, CNN and LC, we use the same architectures used in [9]. We also tested the proposed CorInfoMax-EBD model against the CorInfoMax-EP model of [41]. More details about architectures, implementations, hyperparameter selections, and experimental outputs are provided in the Appendix I.
EBD test accuracy results compared to BP and three error-broadcast methods: DFA without and with entropy regularization (DFA-E) [6], global error vector broadcasting (nonnegative-(NN-GEVB) and mixed-sign-(MS-GEVB))[9] are in Table 1 for both MNIST and CIFAR-10. In addition, the test accuracies for biological CorInfoMax networks trained with EP and EBD methods are in Table 2.
These results show that EBD-trained networks achieve equivalent performance on the MNIST dataset and significantly better performance on the CIFAR-10 dataset. These improvements of EBD in Table 1 over DFA can be attributed to the adaptability of error projection weights in EBD. The improvement of CorInfoMax-EBD over CorInfoMax-EP in Table 2 can be attributed to CorInfoMax-EBD incorporating error decorrelation in updating lateral weights, whereas CorInfoMax-EP relies only on (anti-)Hebbian updates. Particularly noteworthy is the performance of CorInfoMax-EBD, which not only substantially improves upon the original CorInfoMax-EP on CIFAR-10 (e.g., $55.79\%$ vs. $50.97\%$ for $3$-layers with batch size $20$) but also demonstrates encouraging scalability with depth, with a 10-layer CorInfoMax-EBD achieving $96.38\%$ on MNIST and $54.89\%$ on CIFAR-10 using online learning (batch size 1). This highlights EBD’s potential in deeper, more complex biological networks.
@@
(ours) GEVB GEVB (ours)
MNIST MLP $98.1{\pm 0.21}$ $98.2{\pm 0.03}$ $98.1$ $97.7$ $\mathbf{98.7{\pm 0.05}}$ $\underline{98.2{\pm 0.08}}$
CNN $99.1{\pm 0.05}$ $99.1{\pm 0.07}$ $97.7$ $98.2$ $\mathbf{99.5{\pm 0.04}}$ $\underline{99.1{\pm 0.07}}$
LC $\underline{98.9{\pm 0.03}}$ $98.9{\pm 0.04}$ $98.2$ $98.2$ $\mathbf{99.1{\pm 0.04}}$ $\underline{98.9{\pm 0.04}}$
@@ Conclusions.
We introduced the Error Broadcast and Decorrelation algorithm, a biologically plausible alternative to backpropagation. EBD addresses the credit assignment problem by minimizing correlations between layer activations and output errors, offering fresh insights into biologically realistic learning. This approach provides a theoretical foundation for existing error broadcast mechanisms and three-factor learning rules in biological neural networks and facilitates flexible implementations in neuromorphic and artificial neural systems. EBD’s error-broadcasting mechanism aligns with biological processes using local updates, and notably, has proven effective for training deep recurrent biologically-plausible networks (e.g., the 10-layer CorInfoMax-EBD), thereby addressing a key challenge in effectively scaling deep, biologically plausible learning with local rules. Moreover, EBD’s simplicity and parallelism suit efficient hardware, like neuromorphic systems.
@@
The MMSE orthogonality property underlying EBD offers great potential for new algorithms, deeper theoretical understanding, and neural network analysis in both artificial and biological contexts. Further theoretical extensions, drawing from the groundwork laid in Appendix B.2, could focus on deriving tighter convergence guarantees for EBD in practical (finite-width) settings and on investigating the impact of more adaptive choices for the decorrelation functions $g^{(k)}$. In addition, EBD provides theoretical underpinnings for error-broadcast mechanisms with three-factor learning rules, enabling the conversion of two-phase contrastive methods into a single-phase approach. We are currently unaware of similar theoretical properties for alternative loss functions. Finally, our numerical experiments in Appendix J.2 reveal that similar decorrelation behavior occurs for networks trained with backpropagation and categorical cross entropy loss, suggesting that decorrelation may be a general feature of the learning process and an intriguing avenue for further investigation.
@@
Scalability: The current work evaluates EBD on MLP, CNN, LC and recurrent biological networks for image classification tasks like MNIST and CIFAR-10. The 10-layer CorInfoMax-EBD results show promise for scaling EBD with depth in biologically-motivated recurrent architectures using online learning. However, assessing EBD’s performance on significantly larger datasets, or its applicability to diverse large-scale architectures in other domains, remains an important open direction. While related methods like DFA have been explored in such contexts [37], comprehensive empirical validation of EBD itself under those conditions is needed.
@@
6 Acknowledgements
This work was supported by KUIS AI Center Research Award. C.P. is supported by NSF grant DMS-2134157, NSF CAREER Award IIS-2239780, DARPA grant DIAL-FP-038, a Sloan Research Fellowship, and The William F. Milton Fund from Harvard University.
@@ B.2.1 The sufficiency of the nonlinear MMSE orthogonality condition
This theorem provides a strong justification: if we can find an estimator $f_{\Theta}(x)$ whose error $y-f_{\Theta}(x)$ is orthogonal to *all* measurable functions of $x$, then $f_{\Theta}(x)$ is indeed the optimal MMSE estimator. This underpins the EBD framework’s objective.
@@ B.2.2 From orthogonality to hidden units to arbitrary functions: an infinite width perspective
A critical point is that Theorem B.1 requires the estimation error to be orthogonal to *any* measurable function $g(x)$ of the input. However, the EBD algorithm, as practically implemented, enforces orthogonality of the output error $e_{\Theta}(x)=y-f_{\Theta}(x)$ to the activations of the network’s hidden units, $h_{j}^{(k)}(x)$. The question is whether satisfying these more limited orthogonality conditions is sufficient to approach the true MMSE estimator. We argue that in the limit of infinite network width, this can indeed be the case.
The argument relies on the universal approximation capabilities of wide neural networks.
  • Rahimi and Brecht (see also Rahimi and Recht [47]) showed that a hidden layer with random i.i.d. Gaussian weights and appropriately chosen biases (e.g., uniform for Fourier features) can linearly approximate functions within a corresponding Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}_{k}$ associated with a shift-invariant kernel $k$ Rahimi and Recht [48]. Specifically, for a function in $\mathcal{H}_{k}$, $N$ such hidden units (random features) can achieve an expected $L_{2}(P_{X})$ approximation error of order $1/\sqrt{N}$, assuming the input distribution $P_{X}$ has compact support.
  • Building on this, Sun et.al. analyzed random ReLU networks, where hidden units are of the form $\text{ReLU}(w^{\top}x)$ with $w\sim\mathcal{N}(0,I)$ Sun et al. [49]. They demonstrated that the RKHS induced by the corresponding random feature kernel is dense in $L_{2}(P_{X})$ under mild conditions on $P_{X}$. This establishes the universality of random ReLU features for approximating any square-integrable function with the linear combination of that hidden layer units.
Our initial setting for the EBD framework—specifically, a first hidden layer with ReLU activations initialized with random i.i.d. Gaussian weights—is essentially the same as the random ReLU feature setting analyzed in Sun et al. [49]. Although weights change during training, one might hypothesize that in the infinite width limit and with sufficiently controlled learning rates (to limit the deviation of weights from their initial distribution), the set of first hidden layer activations $\{h_{i}^{(1)}(x)\}_{i=1}^{N^{(1)}}$ could retain its universal approximation capability. This is an area for future rigorous analysis.
@@
This theorem is significant because it elucidates the properties of an estimator that achieves the specific orthogonality targeted by the EBD algorithm with respect to its first-layer hidden units. Theorem B.1 established that orthogonality to *all* measurable functions is sufficient for MMSE optimality. Theorem B.2 demonstrates that, *if* an estimator’s error is perfectly orthogonal to its first-layer hidden unit activations, and *if* these activations form a dense basis (a condition motivated by infinite-width random feature networks), then such an estimator indeed converges to the true MMSE solution. This provides a theoretical rationale for the EBD algorithm’s objective: by striving to decorrelate output errors with hidden unit activations, EBD aims to satisfy the premise of this theorem. The result offers a theoretical justification for how achieving orthogonality with respect to a practical, finite set of internal network features can, under idealized conditions of network width and feature richness, lead to overall MMSE optimality. The ”meaningfulness” of these orthogonality constraints, even in high-dimensional spaces, is thus linked to the expressive power of the network’s learned features. The theorem hinges on the denseness of the feature space generated by the first hidden layer and the perfect satisfaction of the orthogonality conditions. Formalizing the conditions under which EBD effectively approximates these conditions and the extent to which universality of features is preserved during training remain important directions for future research.
@@ B.2.3 EBD framework objective
$\langle y-f_{\Theta}(x),h_{j}^{(k)}(x)\rangle=0,\quad\text{for all output components, layers }k,\text{ and units }j.$ (27)
@@ C.1 $\Delta{\mathbf{W}}_{1}$ and $\Delta{\mathbf{b}}_{1}$ calculation
In Section 2.3, we defined the weight update elemet $[\Delta{\mathbf{W}}_{1}]_{ij}$ as follows:
@@ C.2 $\Delta{\mathbf{W}}_{2}$ and $\Delta{\mathbf{b}}_{2}$ calculation
In Section 2.3, we defined the weight update element $[\Delta{\mathbf{W}}_{2}]_{ij}$ involving the derivative of the output error as
@@ C.3 Update corresponding to the layer entropy regularization
[…]
@@ E.2 On gradient truncation and biological plausibility
$\approx\frac{1}{\varepsilon_{k}}\sum_{i=1}^{t}\lambda^{t-i}\|{\mathbf{h}}^{(k+1)}[i]-\mathbf{W}_{ff,*}^{(k)}[m]{\mathbf{h}}^{(k)}[i]\|_{2}^{2}+\varepsilon_{k}\|{\bm{W}}_{ff,*}^{(k)}[m]\|_{F}^{2}+N_{k+1}\log(\varepsilon_{k})$
@@
$\approx\frac{1}{\varepsilon_{k}}\sum_{i=1}^{t}\lambda^{t-i}\|{\mathbf{h}}^{(k)}[i]-\mathbf{W}_{fb,*}^{(k)}[m]{\mathbf{h}}^{(k+1)}[i]\|_{2}^{2}+\varepsilon_{k}\|{\bm{W}}_{fb,*}^{(k)}[m]\|_{F}^{2}+N_{k}\log(\varepsilon_{k}),$
@@ F.3 CorInfoMax-EP
5:Updated weights ${\mathbf{W}}^{(f,k)}[m]$, ${\mathbf{W}}^{(b,k)}[m]$ ,${\mathbf{B}}^{(k)}[m]$
@@
10:$\Delta{\mathbf{W}}^{(f,k)}_{\text{EP}}[m]\leftarrow-\dfrac{\mu^{(d_{f},k)}[m]}{B\beta^{\prime}}\,\left((\overset{\rightarrow}{{\mathbf{E}}}^{(k+1)}[m]{\mathbf{H}}^{(k)}[m]^{T})\Big{|}_{\beta=\beta^{\prime}}-(\overset{\rightarrow}{{\mathbf{E}}}^{(k+1)}[m]{\mathbf{H}}^{(k)}[m]^{T})\Big{|}_{\beta=0}\right)$
11:${\mathbf{W}}^{(f,k)}[m]\leftarrow{\mathbf{W}}^{(f,k)}[m-1]+\Delta{\mathbf{W}}^{(f,k)}_{\text{EP}}[m]$
@@
14:$\Delta{\mathbf{W}}^{(b,k)}_{\text{EP}}[m]\leftarrow-\dfrac{\mu^{(d_{b},k)}[m]}{B\beta^{\prime}}\,\left((\overset{\leftarrow}{{\mathbf{E}}}^{(k)}[m]{\mathbf{H}}^{(k)}[m]^{T})\Big{|}_{\beta=\beta^{\prime}}-(\overset{\leftarrow}{{\mathbf{E}}}^{(k)}[m]{\mathbf{H}}^{(k)}[m]^{T})\Big{|}_{\beta=0}\right)$
15:${\mathbf{W}}^{(b,k)}[m]\leftarrow{\mathbf{W}}^{(b,k)}[m-1]+\Delta{\mathbf{W}}^{(b,k)}_{\text{EP}}[m]$
@@
18:$\Delta{\mathbf{B}}^{(k)}_{\text{E}}[m]\leftarrow-\dfrac{\gamma_{E}}{B}{\mathbf{Z}}^{(k)}[m]{\mathbf{Z}}^{(k)}[m]^{T}$
19:${\mathbf{B}}^{(k)}[m]\leftarrow\frac{1}{\lambda_{E}}{\mathbf{B}}^{(k)}[m]+\Delta{\mathbf{B}}^{(k)}_{\text{E}}[m]$
@@ G.1.2 Additional cost of Entropy Regularization
[…]
@@
[…]
@@
[…]

### v3

@@ 4 Numerical experiments
In this section, we evaluate the performance of the proposed Error Broadcast and Decorrelation (EBD) approach on benchmark datasets: MNIST [48] and CIFAR-10/100 [49]. For experiments with MNIST and CIFAR-10 involving MLP, CNN and LC, we use the same architectures used in [9]; while for CIFAR-100 we adopt a CNN architecture closely following that of [41]. We also tested the proposed CorInfoMax-EBD model against the CorInfoMax-EP model of [45]. More details about architectures, implementations, hyperparameter selections, and experimental outputs are provided in the Appendix I.
EBD test accuracy results compared to BP (with MSE criterion) and three error-broadcast methods: DFA without and with entropy regularization (DFA-E) [6], global error vector broadcasting (nonnegative-(NN-GEVB) and mixed-sign-(MS-GEVB))[9] are in Table 1 for both MNIST and CIFAR-10. Under our training setup, BP yielded comparable test accuracies for these datasets with both MSE and Cross-Entropy losses, though we report only the MSE results. In addition, CIFAR-100 results of EBD compared to BP (with Cross-Entropy criterion) and DFA is given in Table 2. Lastly, the test accuracies for biological CorInfoMax networks trained with EP and EBD methods are in Table 3.
These results show that EBD-trained networks achieve equivalent performance on the MNIST dataset and significantly better performance on the CIFAR-10 and CIFAR-100 datasets compared to other error broadcasting methods. These improvements of EBD in Table 1 over DFA can be attributed to the adaptability of error projection weights in EBD. The improvement of CorInfoMax-EBD over CorInfoMax-EP in Table 3 can be attributed to CorInfoMax-EBD incorporating error decorrelation in updating lateral weights, whereas CorInfoMax-EP relies only on (anti-)Hebbian updates. Particularly noteworthy is the performance of CorInfoMax-EBD, which not only substantially improves upon the original CorInfoMax-EP on CIFAR-10 (e.g., $55.79\%$ vs. $50.97\%$ for $3$-layers with batch size $20$) but also demonstrates encouraging scalability with depth, with a 10-layer CorInfoMax-EBD achieving $96.38\%$ on MNIST and $54.89\%$ on CIFAR-10 using online learning (batch size 1). This highlights EBD’s potential in deeper, more complex biological networks.
@@
(ours) GEVB GEVB (MSE) (ours)
MNIST MLP $98.1{\pm 0.21}$ $\underline{98.2{\pm 0.03}}$ $98.1$ $97.7$ $\mathbf{98.7{\pm 0.05}}$ $\underline{98.2{\pm 0.08}}$
CNN $\underline{99.1{\pm 0.05}}$ $\underline{99.1{\pm 0.07}}$ $97.7$ $98.2$ $\mathbf{99.5{\pm 0.04}}$ $\underline{99.1{\pm 0.07}}$
LC $\underline{98.9{\pm 0.03}}$ $\underline{98.9{\pm 0.04}}$ $98.2$ $98.2$ $\mathbf{99.1{\pm 0.04}}$ $\underline{98.9{\pm 0.04}}$
@@
Table 2: Accuracy (%) results for the CNN on CIFAR-100.
Dataset Model DFA BP (CE) EBD (ours)
CIFAR-100 CNN $41.9{\pm 0.32}$ $\mathbf{60.5{\pm 0.17}}$ $\underline{45.9{\pm 0.69}}$
@@ Conclusions.
We introduced the Error Broadcast and Decorrelation framework, a biologically plausible alternative to backpropagation. EBD addresses the credit assignment problem by minimizing correlations between layer activations and output errors, offering fresh insights into biologically realistic learning. This approach provides a theoretical foundation for existing error broadcast mechanisms and three-factor learning rules in biological neural networks and facilitates flexible implementations in neuromorphic and artificial neural systems. EBD’s error-broadcasting mechanism aligns with biological processes using local updates, and notably, has proven effective for training deep recurrent biologically-plausible networks (e.g., the 10-layer CorInfoMax-EBD), thereby addressing a key challenge in effectively scaling deep, biologically plausible learning with local rules. Moreover, EBD’s simplicity and parallelism suit efficient hardware, like neuromorphic systems.
@@
The MMSE orthogonality property underlying EBD offers significant promise for new algorithms, deeper theoretical understanding, and neural network analysis in both artificial and biological contexts. Further theoretical extensions, drawing from the groundwork laid in Appendix B.2, could focus on deriving tighter convergence guarantees for EBD in practical (finite-width) settings and on investigating the impact of more adaptive choices for the decorrelation functions $g^{(k)}$. In addition, EBD provides theoretical underpinnings for error-broadcast mechanisms with three-factor learning rules, enabling the conversion of two-phase contrastive methods into a single-phase approach. We are currently unaware of similar theoretical properties for alternative loss functions. Finally, our numerical experiments in Appendix J.2 reveal that similar decorrelation behavior occurs for networks trained with backpropagation and categorical cross entropy loss, suggesting that decorrelation may be a general feature of the learning process and an intriguing avenue for further investigation.
@@
Scalability: The current work evaluates EBD on MLP, CNN, LC and recurrent biological networks for image classification tasks like MNIST and CIFAR-10. Results on the 10-layer CorInfoMax-EBD demonstrate the potential to scale EBD to deeper, biologically realistic recurrent architectures using online learning. However, assessing EBD’s performance on significantly larger datasets, or its applicability to diverse large-scale architectures in other domains, remains an important open direction. While related methods like DFA have been explored in such contexts [41], comprehensive empirical validation of EBD itself under those conditions is needed.
@@
Acknowledgements
This work was supported by KUIS AI Center Research Award. C.P. was supported by an NSF CAREER Award (IIS-2239780) and a Sloan Research Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
@@ B.2.1 The sufficiency of the nonlinear MMSE orthogonality condition
This theorem provides a strong justification: if we can find an estimator $f_{\Theta}(x)$ whose error $y-f_{\Theta}(x)$ is orthogonal to all measurable functions of $x$, then $f_{\Theta}(x)$ is indeed the optimal MMSE estimator. This underpins the EBD framework’s objective.
@@ B.2.2 From orthogonality to hidden units to arbitrary functions: an infinite width perspective
A critical point is that Theorem B.1 requires the estimation error to be orthogonal to any measurable function $g(x)$ of the input. However, the EBD algorithm, as practically implemented, enforces orthogonality of the output error $e_{\Theta}(x)=y-f_{\Theta}(x)$ to the activations of the network’s hidden units, $h_{j}^{(k)}(x)$. The question is whether satisfying these more limited orthogonality conditions is sufficient to approach the true MMSE estimator. We argue that in the limit of infinite network width, this can indeed be the case.
The argument relies on the universal approximation capabilities of wide neural networks:
  • Rahimi and Recht showed that a hidden layer with random i.i.d. Gaussian weights and appropriately chosen biases (e.g., uniform for Fourier features) can linearly approximate functions within a corresponding Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}_{k}$ associated with a shift-invariant kernel $k$ [51, 52]. Specifically, for a function in $\mathcal{H}_{k}$, $N$ such hidden units (random features) can achieve an expected $L_{2}(P_{X})$ approximation error of order $1/\sqrt{N}$, assuming the input distribution $P_{X}$ has compact support.
  • Building on this, Sun et.al. analyzed random ReLU networks, where hidden units are of the form $\text{ReLU}(w^{\top}x)$ with $w\sim\mathcal{N}(0,I)$ [53]. They demonstrated that the RKHS induced by the corresponding random feature kernel is dense in $L_{2}(P_{X})$ under mild conditions on $P_{X}$. This establishes the universality of random ReLU features for approximating any square-integrable function with the linear combination of that hidden layer units.
Our initial setting for the EBD framework—specifically, a first hidden layer with ReLU activations initialized with random i.i.d. Gaussian weights—is essentially the same as the random ReLU feature setting analyzed in [53]. Although weights change during training, one might hypothesize that in the infinite width limit and with sufficiently controlled learning rates (to limit the deviation of weights from their initial distribution), the set of first hidden layer activations $\{h_{i}^{(1)}(x)\}_{i=1}^{N^{(1)}}$ could retain its universal approximation capability. This is an area for future rigorous analysis.
@@
This theorem is significant because it elucidates the properties of an estimator that achieves the specific orthogonality targeted by the EBD algorithm with respect to its first-layer hidden units. Theorem B.1 established that orthogonality to all measurable functions is sufficient for MMSE optimality. Theorem B.2 demonstrates that, if an estimator’s error is perfectly orthogonal to its first-layer hidden unit activations, and if these activations form a dense basis (a condition motivated by infinite-width random feature networks), then such an estimator indeed converges to the true MMSE solution. This provides a theoretical rationale for the EBD algorithm’s objective: by striving to decorrelate output errors with hidden unit activations, EBD aims to satisfy the premise of this theorem. The result offers a theoretical justification for how achieving orthogonality with respect to a practical, finite set of internal network features can, under idealized conditions of network width and feature richness, lead to overall MMSE optimality. The ”meaningfulness” of these orthogonality constraints, even in high-dimensional spaces, is thus linked to the expressive power of the network’s learned features. The theorem hinges on the denseness of the feature space generated by the first hidden layer and the perfect satisfaction of the orthogonality conditions. Formalizing the conditions under which EBD effectively approximates these conditions and the extent to which universality of features is preserved during training remain important directions for future research.
@@ B.2.3 EBD framework objective
[…]
@@ C.1 $\Delta{\mathbf{W}}_{1}$ and $\Delta{\mathbf{b}}_{1}$ calculation
In Section 2.3, we defined the weight update elemet $\Delta{\mathbf{W}}_{1}]_{ij}$ as follows:
@@ C.2 $\Delta{\mathbf{W}}_{2}$ and $\Delta{\mathbf{b}}_{2}$ calculation
In Section 2.3, we defined the weight update element $\Delta{\mathbf{W}}_{2}]_{ij}$ involving the derivative of the output error as
@@ C.3 Update corresponding to the layer entropy regularization
${\mathbf{R}}^{(k)}_{\mathbf{h}}[m]$ $=\lambda_{E}{\mathbf{R}}^{(k)}_{\mathbf{h}}[m-1]+(1-\lambda_{E})\frac{1}{B}{{\mathbf{H}}^{(k)}[m]}{{\mathbf{H}}^{(k)}[m]^{T}},\hskip 14.45377pt\text{and,}$
@@ E.2 On gradient truncation and biological plausibility
$\hskip-122.85876pt\approx\frac{1}{\varepsilon_{k}}\sum_{i=1}^{t}\lambda^{t-i}\|{\mathbf{h}}^{(k+1)}[i]-\mathbf{W}_{ff,*}^{(k)}[m]{\mathbf{h}}^{(k)}[i]\|_{2}^{2}+\varepsilon_{k}\|{\bm{W}}_{ff,*}^{(k)}[m]\|_{F}^{2}+N_{k+1}\log(\varepsilon_{k})$
@@
$\hskip-122.85876pt\approx\frac{1}{\varepsilon_{k}}\sum_{i=1}^{t}\lambda^{t-i}\|{\mathbf{h}}^{(k)}[i]-\mathbf{W}_{fb,*}^{(k)}[m]{\mathbf{h}}^{(k+1)}[i]\|_{2}^{2}+\varepsilon_{k}\|{\bm{W}}_{fb,*}^{(k)}[m]\|_{F}^{2}+N_{k}\log(\varepsilon_{k}),$
@@ F.3 CorInfoMax-EP
5:Updated weights ${\mathbf{W}}^{(f,k)}[m]$, ${\mathbf{W}}^{(b,k)}[m]$ ${\mathbf{B}}^{(k)}[m]$
@@
[…]
@@
[…]
@@
[…]
@@ G.1.2 Additional cost of Entropy Regularization
    $\nabla_{\mathbf{W}^{(k)}}\,J^{(k)}[m]\;=\;2\,\frac{1-\lambda_{E}}{B}\;\Bigl[(\mathbf{R}_{h}^{(k)}[m]+\epsilon\,\mathbf{I})^{-1}\,\mathbf{H}^{(k)}[m]\;\odot\;f^{\prime}\!\bigl(\mathbf{W}^{(k)}\,\mathbf{H}^{(k-1)}[m]+\mathbf{b}^{(k)}\,\mathbf{1}^{T}\bigr)\Bigr]\;\bigl(\mathbf{H}^{(k-1)}[m]\bigr)^{T},$
@@
        The matrix–matrix product $\mathbf{H}^{(k)}[m]\in\mathbb{R}^{N^{(k)}\times B}$ times its transpose in $\mathbb{R}^{B\times N^{(k)}}$ yields an $N^{(k)}\times N^{(k)}$ matrix, costing $\mathcal{O}\bigl(N^{(k)\,2}\,B\bigr)$.
@@
    Hence, using the Woodbury identity is beneficial whenever $B$ is much smaller than $N^{(k)}$, because $N^{(k)\,2}\,B+B^{3}\ll(N^{(k)})^{3}$.